OpenStudy (anonymous):

Find the area of the region bounded by the curve of x=y^2-1 and the y-axis.

OpenStudy (nincompoop):

how are you going to solve this?

OpenStudy (anonymous):

By hand without graphing.

OpenStudy (nincompoop):

and how are you going to do that exactly?

OpenStudy (anonymous):

I don't know, that's why I asked the question.

OpenStudy (nincompoop):

so you're doing calculus now?

OpenStudy (anonymous):

Yes.

OpenStudy (nincompoop):

what have you learned in the past about area problems?

OpenStudy (anonymous):

Take the integral after finding the limits of integration.

OpenStudy (nincompoop):

okay, go ahead

OpenStudy (nincompoop):

set up the integral notation

OpenStudy (anonymous):

But should the integrand be y^2-1? Or y=sqrt(x+1)?

OpenStudy (nincompoop):

well does it matter?

OpenStudy (nincompoop):

would it satisfy the equation if the integral is set up like this? \[\int\limits (y^2-1) dy \]

OpenStudy (anonymous):

I think so. But don't I need to find the limits of integration by setting y^2-1=0 and solve for y?

OpenStudy (nincompoop):

and I don't know why you mention about limits of integration when the integration problem is indefinite

OpenStudy (anonymous):

Are you sure? Because this is about finding the area. The integral must be definite.

OpenStudy (nincompoop):

so how are you going to find the limits?

OpenStudy (nincompoop):

from what to what?

OpenStudy (anonymous):

That's the part I don't get. If I set y^2-1=0 and solve for y, I get 1 and -1. That means I need to take the integral from -1 to 1. And I was asking you if that's the right way to do it.

OpenStudy (nincompoop):

okay then set it up that way, but when you do that what happens to your x?

OpenStudy (nincompoop):

can you make sure first that you've copied down the problem correctly?

OpenStudy (nincompoop):